OpenStudy (anonymous):

Eliminate t from the following parametric equations to find there equivalent Cartesian forms x=2t cos(t), y=3t sin(t), z = t Im not sure where to start. If that other multiple of t was not there I would take the the inverse trig function to solve for t. Need a shove in the right direction.

OpenStudy (asnaseer):

ok, first replace t with z in the equations for x and y

OpenStudy (asnaseer):

then get and expression for cos(z) and sin(z)

OpenStudy (asnaseer):

finally use the rule:\[\cos^2(z)+\sin^2(z)=1\]to get rid of the cos/sin.

OpenStudy (anonymous):

ok..ill try to work through it

OpenStudy (asnaseer):

let me know what you find and I'll check it for you.

OpenStudy (anonymous):

thank you , ill bb in a min

OpenStudy (anonymous):

Ok, im stumped. If I replace z with t I get x = 2z cos(z) y=3z sin(z) if I get expressoins for cos and sin x/2z = cos(z) and y/3z = sin(z) Im just now sure what good that does?

OpenStudy (asnaseer):

use the rule I showed you above, i.e.:\[\cos^2(z)+\sin^2(z)=1\]